OpenStudy (anonymous):

A box of staples has a length of 6 cm, a width of 7 cm, and a volume of 378 cm cubed. What is the height of the box?

OpenStudy (anonymous):

V of rectangle = l * w * h V=378cm 378cm = 6cm * 7 cm * h 378cm =42 * h h= 9 cm